# Tracking objects that change in appearance with phase synchrony

### v1

@@ Abstract
Objects we encounter often change appearance as we interact with them. Changes in illumination (shadows), object pose, or movement of nonrigid objects can drastically alter available image features. How do biological visual systems track objects as they change? It may involve specific attentional mechanisms for reasoning about the locations of objects independently of their appearances — a capability that prominent neuroscientific theories have associated with computing through neural synchrony. We computationally test the hypothesis that the implementation of visual attention through neural synchrony underlies the ability of biological visual systems to track objects that change in appearance over time. We first introduce a novel deep learning circuit that can learn to precisely control attention to features separately from their location in the world through neural synchrony: the complex-valued recurrent neural network (CV-RNN). Next, we compare object tracking in humans, the CV-RNN, and other deep neural networks (DNNs), using FeatureTracker: a large-scale challenge that asks observers to track objects as their locations and appearances change in precisely controlled ways. While humans effortlessly solved FeatureTracker, state-of-the-art DNNs did not. In contrast, our CV-RNN behaved similarly to humans on the challenge, providing a computational proof-of-concept for the role of phase synchronization as a neural substrate for tracking appearance-morphing objects as they move about.
@@ 1 Introduction
The neural substrates for tracking objects by re-recognition or motion integration have been the focus of extensive studies over the past half-century, and the current consensus is that distinct neural circuits are responsible for each strategy (Lettvin et al., 1959; Takemura et al., 2013; Kim et al., 2014; Adelson & Bergen, 1985; Frye, 2015; Pylyshyn, 2006). Much less progress has been made in characterizing how visual systems track objects as their appearances change (Fig. 1c,d), although visual attention likely plays a critical role in tracking  (Blaser et al., 2000). Indeed, visual attention is considered essential for solving many visual challenges that occur during object tracking, such as maintaining the location of an object even as it is occluded from view (Koch & Ullman, 1987; Roelfsema et al., 1998; Busch & VanRullen, 2010; Herrmann & Knight, 2001; Pylyshyn & Storm, 1988; Pylyshyn, 2006). We hypothesize that visual attention similarly helps when tracking objects that change appearance by maintaining information about their location in the world independently from their appearances.
@@
Figure 1: How do Biological visual systems track the object tagged by the yellow arrow? (a) Sometimes, the object’s appearance makes it easy to track (Pylyshyn, 2006; Pylyshyn & Storm, 1988). (b) Other times, when objects look similar, the target can be tracked by following its motion through the world (Lettvin et al., 1959; Takemura et al., 2013; Kim et al., 2014; Adelson & Bergen, 1985; Frye, 2015; Linsley et al., 2021). Here, we investigate a computational problem that has received far less attention: how do biological visual systems track objects when their colors, textures (c), or shapes (d) change over time? (e) We developed the FeatureTracker challenge to systematically evaluate humans and machine vision systems on this problem. In FeatureTracker, observers watch videos containing objects that change in color and/or shape over time, and have to decide if the target object, which begins in the red square (circled in white for clarity), winds up in the blue square by the end of a video. When presented with a FeatureTracker video, one possible strategy suggested by neuroscience theories is that oscillatory activity of neural populations can keep track of different objects over time. Specifically, the target is encoded by a population of neurons that fire with a timing that differs from that of the population that responds to the distractors. We approximate the cycle of the oscillation with complex-valued neurons. In the CV-RNN, the phase of a complex-valued neuron represents the object encoded by this neuron. The CV-RNN thus learns to tag the target with a phase value different from the phase value of the distractors.
@@
Recent work proposed using complex-valued representations to implement neural synchrony in artificial models (Reichert & Serre, 2013; Löwe et al., 2022; Stanić et al., 2023). According to the framework proposed by Reichert & Serre (2013), each neuron in an artificial neural network can be represented as a complex number where the magnitude encodes for specific object features, and the phase groups the features of different objects. Such representations allow the modeling of various neuroscience theories (Singer & Gray, 2003; Singer, 2007; 2009) related to the role of neural synchrony. Here, we investigate whether the use of complex-valued representations to implement neural synchrony can help to solve the FeatureTracker challenge through large-scale computational experiments (see Fig. 1e).
@@
  • Humans are exceptionally accurate at tracking objects in the FeatureTracker challenge as they move through the world and change in color, shape, or color and shape.
@@
  • Inspired by theory on how populations of neurons implement solutions to the binding problems of FeatureTracker, we incorporated a novel mechanism for computing attention through neural synchrony, using complex-valued representations, in a recurrent neural network architecture, which we call the complex-valued recurrent neural network (CV-RNN). The CV-RNN approaches human performance and decision-making on FeatureTracker.
  • Our findings establish proof-of-concept that neural synchrony may support object tracking in humans, and can induce similar capabilities in artificial visual systems. We release FeatureTracker data, code, and human psychophysics at https://anonymous.4open.science/r/feature_tracker-2CA3 to help the field investigate this gap between human and machine vision.
@@ 2 Background and related work
The empirical finding that alpha/beta (12–30Hz) and gamma (>30Hz) oscillations tend to be anti-correlated in primate cortex has motivated the development of theories on how the temporal synchronization of different groups of neurons may reflect an overarching computational strategy of brains. In the communication-through-coherence (CTC) theory, Fries (2015) proposed that alpha/beta activity carries top-down attentional signals, which reflect information about the current context and goals. Others have expanded on this theory to propose that these top-down signals can be spatially localized in the cortex to multiplex attentional computations independently of the features encoded by neurons (Miller et al., 2024). While there have been many different theories proposed on how computing through oscillations works (McLelland & VanRullen, 2016; Lisman & Jensen, 2013; Grossberg, 1976; Milner, 1974; Mioche & Singer, 1989), here we assume an induced oscillation and study synchrony as a mechanism for visual routines and its potential for implementing object tracking in brains.
@@
Figure 2: Neural synchrony helps track objects that change in appearance. (a) The shell game is designed to probe how a neural network, with the functional constraints of biological visual systems, could track objects as they change in appearance between frames one and two. Are the two images the same, or has the objects’ color and/or orientation flipped (three possible responses)? (b) We tested a simplified model of the hierarchical visual system on the task, which consisted of two layers of neurons: (i) a convolutional layer with high-resolution feature maps, followed by (ii) a spatial average pooling of neuron responses and a layer of recurrently connected neurons (McLelland & VanRullen, 2016). 1c/2c are object colors, 1o/2o are object orientations; the loss of spatial resolution between the layers caused these object features to interfere with each other. The model is able to detect the features present in the frame (red and blue color, as well as square and diamond orientations), but fails at binding the color and orientation with the position – hence cannot differentiate Frame 1 from Frame 2. (c, d) The same architecture can learn to solve the task with a complex-valued mechanism for neural synchrony, in which the magnitude of neurons captures object appearances, and the phase captures object locations.
Generalization and shortcut learning in deep neural networks
A drawback of deep neural networks’ (DNNs) great power is their tendency to learn spurious correlations between inputs and labels, which can lead to poor generalization (Barbu et al., 2019; Geirhos et al., 2020b). Moreover, while object classification models have grown more accurate over the past decade — now matching and sometimes exceeding human performance (Shankar et al., 2020) — they have done so by learning recognition strategies that are becoming progressively less aligned with humans (Fel* et al., 2022; Linsley et al., 2023b; a). Synthetic datasets like FeatureTracker are useful for understanding why this misalignment occurs and guiding the development of novel architectures that can address it. For example, visual reasoning differences between humans and DNNs have been probed by challenges like Bongard-LOGO (Nie et al., 2020), cABC (Kim* et al., 2020), SVRT (Fleuret et al., 2011), and PSVRT (Kim et al., 2018). The PathFinder challenge, which was originally developed to investigate the ability of observers to trace long curves in clutter (Linsley et al., 2018), was used to optimize Transformer and modern state space model architectures (Tay et al., 2021; Gu et al., 2021; Smith et al., 2022). The most similar challenge to our FeatureTracker is PathTracker, which tested whether observers could track one object in a swarm of identical-looking objects as they briefly occlude each other while they move around (Linsley et al., 2021). Here, we extend PathTracker by adding parametric control over the shape and color of objects to test tracking as object appearances smoothly change.
@@ 4 The FeatureTracker challenge
Our shell game was a highly simplified test of object tracking. In the real world, objects and their appearance and spatial features evolve smoothly and predictably over time. To better understand the capabilities of our CV-RNN — and neural synchrony — to track objects under such conditions, we next developed the FeatureTracker challenge. In FeatureTracker, observers watch a video and decide if a target object, which begins in a red square, travels to a blue square by the end of the video as opposed to a distractor (Fig. 4). This task is challenging for two reasons: (i) each video contains distractor objects that sometimes pass by and occlude the target object, and (ii) the color, shape, or color and shape of all objects in each video morph over time in precisely controlled ways.
@@
We created the FeatureTracker challenge to probe how well observers could track objects as their appearances changed in familiar or unfamiliar ways. We did this by sampling the starting state of each object’s color and/or shape from distributions that we varied from training to test time. To elaborate, object appearances for training were sampled from one distribution (Fig. 4, red cube); then observers were tested on objects with appearances sampled in four different ways: (i) colors/shapes from the same distribution, (ii) colors from a different distribution but shapes from the same distribution, (iii) colors from the same distribution but shapes from a different distribution, or (iv) colors from a different distribution and shapes from a different distribution (Fig. 4 and SI A.7.2). We systematically evaluated the abilities of humans and machine vision systems to track objects with changing appearances by comparing their performances and decision strategies on each of these test sets.
@@ Results
Figure 6: Neural synchrony causes the CV-RNN to act more like humans than any other model tested. Decision correlations of humans and models for each testing condition: features out-of-distribution conditions (a) and object evolution conditions (b), computed using the Error consistency measure (Geirhos et al., 2020b). We present the same videos to humans and models and compare the errors made by humans and models. Each dot represents the error consistency averaged across human subjects. A higher number represents similar errors between humans and the model. The grey box represents the human inter-subject agreement. (c) Visualization of the phases of the hidden state of the complex attention mechanisms of our CV-RNN. The model affects a phase value for the target which is different from the one of the distractors and the background, and remains consistent across frames.
@@
All of the models that we tested, with the exception of the two based on visual Transformers, rivaled or exceeded human performance on videos with objects that were sampled from the same distribution as models and humans were trained on (Fig. 5a, top-left). The models performed similarly when tested on videos of objects with similar colors as training but different shapes (Fig. 5a, bottom-left). However, model performance fell precipitously when object colors were sampled from a different distribution at test time (Fig. 5a, right). In both of these conditions (same shapes/different colors, different shapes/different colors), the CV-RNN performed significantly better than the other models, which were close to chance accuracy.
To further probe the abilities of models and humans to solve FeatureTracker, we generated versions where either the shape, color, or shape and color of objects were fixed to a single value. We first calibrated the performance of models and a new set of human participants against the first experiment by testing both on FeatureTracker videos where the color and shapes of objects varied according to the training distribution. In this case, humans were only rivaled by the CV-RNN, InT RNN, and two of the 3D-CNNs (Fig. 5b, top-left). Only the CV-RNN’s performance fell within the human confidence interval on the remaining conditions, which consisted of objects with fixed shapes but varying colors (Fig. 5b, bottom-left), varying shapes but fixed colors (top-right), or fixed colors and shapes (bottom-right; see SI Figs. A.12 and  A.13 for extended benchmarks). We also include two additional experiments controling the ability of the models to generalize to in-distribution features but out-of-distribution feature trajectories or to occlusions of the target in Figs. A.14 and A.15.
@@
To better understand how well our implementation of neural synchrony captures human strategies for object tracking, we next compared the errors it makes on FeatureTracker to humans. To test this, we turned to the method introduced by Geirhos et al. (2020a), which uses Cohen’s $\kappa$ to compute error correlations between observers while controlling for their task accuracy. For any given version of FeatureTracker, we then computed human-to-human (intersubject) error consistency as the average $\kappa$ between all combinations of humans and model-to-human error consistency as the average $\kappa$ between a model and all humans.
CV-RNN errors were far more consistent with humans than other models (Fig. 6a). The CV-RNN’s error consistency fell within the human intersubject interval on 6/8 versions of FeatureTracker, and was most similar to humans in the remaining two versions of FeatureTracker (expanded results in SI A.13.2, Figs.A.16 and A.17). The InT RNN was more consistent than the CV-RNN on FeatureTracker videos that were sampled from the same distribution as models were trained on (Fig. 6a, top-left), but was otherwise far less aligned with humans than the CV-RNN. In contrast, the remaining DNNs tended to make errors on different videos than humans.
The CV-RNN’s ability to learn neural synchrony makes it more similar to human accuracy and decision-making on FeatureTracker than any other model we tested. To understand how the CV-RNN uses neural synchrony, we visualized the phase of the model’s complex-valued hidden state $\phi_{xy}$ over FeatureTracker videos. The model learned to use the phase of its hidden state to track the position of targets independently of their color or shape. When the model correctly classified FeatureTracker videos, the phase of its attention smoothly tracked the target object and inhibited occluding distractors. When the model was incorrect, its phase tag jumped from object to object as it searched for the target (see SI A.8.3 for more examples).

### v3

@@ Abstract
Objects we encounter often change appearance as we interact with them. Changes in illumination (shadows), object pose, or the movement of non-rigid objects can drastically alter available image features. How do biological visual systems track objects as they change? One plausible mechanism involves attentional mechanisms for reasoning about the locations of objects independently of their appearances — a capability that prominent neuroscience theories have associated with computing through neural synchrony. Here, we describe a novel deep learning circuit that can learn to precisely control attention to features separately from their location in the world through neural synchrony: the complex-valued recurrent neural network (CV-RNN). Next, we compare object tracking in humans, the CV-RNN, and other deep neural networks (DNNs), using FeatureTracker: a large-scale challenge that asks observers to track objects as their locations and appearances change in precisely controlled ways. While humans effortlessly solved FeatureTracker, state-of-the-art DNNs did not. In contrast, our CV-RNN behaved similarly to humans on the challenge, providing a computational proof-of-concept for the role of phase synchronization as a neural substrate for tracking appearance-morphing objects as they move about.
@@ 1 Introduction
The neural substrates for tracking objects by re-recognition or motion integration have been the focus of extensive studies over the past half-century. The current consensus is that distinct neural circuits are responsible for each strategy (Lettvin et al., 1959; Takemura et al., 2013; Kim et al., 2014; Adelson & Bergen, 1985; Frye, 2015; Pylyshyn, 2006). Much less progress has been made in characterizing how visual systems track objects as their appearances change (Fig. 1c,d). However, visual attention likely plays a critical role in tracking  (Blaser et al., 2000). Visual attention is considered essential to solve many visual challenges that occur during object tracking, such as maintaining the location of an object even as it is occluded from view (Koch & Ullman, 1987; Roelfsema et al., 1998; Busch & VanRullen, 2010; Herrmann & Knight, 2001; Pylyshyn & Storm, 1988; Pylyshyn, 2006). We hypothesize that visual attention similarly helps when tracking objects that change appearance by maintaining information about their location in the world independently of their appearances.
@@
Figure 1: How do Biological visual systems track the object tagged by the yellow arrow? (a) Sometimes, the object’s appearance makes it easy to track (Pylyshyn, 2006; Pylyshyn & Storm, 1988). (b) Other times, when objects look similar, the target can be tracked by following its motion through the world (Lettvin et al., 1959; Takemura et al., 2013; Kim et al., 2014; Adelson & Bergen, 1985; Frye, 2015; Linsley et al., 2021). Here, we investigate a computational problem that has received far less attention: how do biological visual systems track objects when their colors, textures (c), or shapes (d) change over time? (e) We developed the FeatureTracker challenge to systematically evaluate humans and machine vision systems on this problem. In FeatureTracker, observers watch videos containing objects that change in color and/or shape over time, and have to decide if the target object, which begins in the red square (circled in white for clarity), ends up in the blue square by the end of a video. When presented with a FeatureTracker video, one possible strategy suggested by neuroscience theories is that the oscillatory activity of neural populations can keep track of different objects over time. Specifically, the target is encoded by a population of neurons that fire with a timing that differs from that of the population that responds to the distractors Astrand et al. (2020). We approximate the cycle of the oscillation with complex-valued neurons. In the CV-RNN, the phase of a complex-valued neuron represents the object encoded by this neuron. The CV-RNN thus learns to tag the target with a phase value different from the phase value of the distractors.
@@
Previous work proposed using complex-valued representations in RNNs (Lee et al., 2022), and in other architectures to implement neural synchrony in artificial models (Reichert & Serre, 2013; Löwe et al., 2022; Stanić et al., 2023). According to the framework proposed by Reichert & Serre (2013), each neuron in an artificial neural network can be represented as a complex number where the magnitude encodes for specific object features, and the phase groups the features of different objects. Such representations allow the modeling of various neuroscience theories (Singer & Gray, 2003; Singer, 2007; 2009) related to the role of neural synchrony. Here, we investigate whether the use of complex-valued representations to implement neural synchrony can help to solve the FeatureTracker challenge through large-scale computational experiments (see Fig. 1e).
@@
  • Humans are exceptionally accurate at tracking objects in the FeatureTracker challenge as these objects move through the world and change in color, shape, or both.
@@
  • Inspired by neuroscience theories on how populations of neurons implement solutions to the binding problem of FeatureTracker, we incorporated a novel mechanism for computing attention through neural synchrony, using complex-valued representations, in a recurrent neural network architecture, which we call the complex-valued recurrent neural network (CV-RNN). The CV-RNN approaches human performance and decision-making on FeatureTracker.
  • Our findings establish a proof-of-concept that neural synchrony may support object tracking in humans, and can induce similar capabilities in artificial visual systems. We release FeatureTracker data, code, and human psychophysics at https://github.com/S4b1n3/feature_tracker to help the field investigate this gap between human and machine vision.
@@ 2 Background and related work
The empirical finding that alpha/beta (12–30Hz) and gamma (>30Hz) oscillations tend to be anti-correlated in primate cortex has motivated the development of theories on how the temporal synchronization of different groups of neurons may reflect an overarching computational strategy of brains. In the communication-through-coherence (CTC) theory, Fries (2015) proposed that alpha/beta activity carries top-down attentional signals, which reflect information about the current context and goals. Others have expanded on this theory to suggest that these top-down signals can be spatially localized in the cortex to multiplex attentional computations independently of the features encoded by neurons (Miller et al., 2024). While there have been many different theories proposed on how computing through oscillations works (McLelland & VanRullen, 2016; Lisman & Jensen, 2013; Grossberg, 1976; Milner, 1974; Mioche & Singer, 1989), here we assume an induced oscillation and study synchrony as a mechanism for visual routines and its potential for implementing object tracking in brains.
@@
Figure 2: Neural synchrony helps track objects that change in appearance. (a) The shell game is designed to probe how a neural network, with the functional constraints of biological visual systems, could track objects as they change in appearance between frames one and two. Are the two images the same, or has the objects’ color and/or orientation flipped (three possible responses)? (b) We tested a simplified model of the hierarchical visual system on the task, which consisted of two layers of neurons: (i) a convolutional layer with high-resolution feature maps, followed by (ii) a spatial average pooling of neuron responses and a layer of recurrently connected neurons (McLelland & VanRullen, 2016). 1c/2c are object colors, 1o/2o are object orientations; the loss of spatial resolution between the layers causes these object features to interfere. The model can detect the features present in the frame (red and blue color, as well as square and diamond orientations), but fails at binding the color and orientation with the position – hence cannot differentiate Frame 1 from Frame 2. (c, d) The same architecture can learn to solve the task with a complex-valued mechanism for neural synchrony, in which the magnitude of neurons captures object appearances, and the phase captures object locations.
Generalization and shortcut learning in DNNs
A drawback of DNNs’ great power is their tendency to learn spurious correlations between inputs and labels, which can lead to poor generalization (Barbu et al., 2019; Geirhos et al., 2020b). Moreover, while object classification models have grown more accurate over the past decade — now matching and sometimes exceeding human performance (Shankar et al., 2020) — they have done so by learning recognition strategies that are becoming progressively less aligned with humans (Fel* et al., 2022; Linsley et al., 2023b; a). Synthetic datasets like FeatureTracker are useful for understanding why this misalignment occurs and guiding the development of novel architectures that can address it. The PathFinder challenge, which was originally developed to investigate the ability of observers to trace long curves in clutter (Linsley et al., 2018), was used to optimize Transformer and modern state space model architectures (Tay et al., 2021; Gu et al., 2021; Smith et al., 2022). The most similar challenge to our FeatureTracker is PathTracker, which tested whether observers could track one object in a swarm of identical-looking objects as they briefly occlude each other while they move around (Linsley et al., 2021). Here, we extend PathTracker by adding parametric control over the shape and color of objects to test tracking as object appearances smoothly change.
@@ 4 The FeatureTracker challenge
Our shell game was a highly simplified test of object tracking. In the real world, objects and their appearance and spatial features evolve smoothly and predictably over time. To better understand the capabilities of our CV-RNN — and neural synchrony — to track objects under such conditions, we next developed the FeatureTracker challenge. In FeatureTracker, observers watch a video and decide if a target object, which begins in a red square, travels to a blue square by the end of the video as opposed to a distractor (Fig. 4). This task is challenging for two reasons: (i) each video contains distractor objects that sometimes pass by and occlude the target object, and (ii) the color and/or shape of all objects in each video morph over time in precisely controlled ways.
@@
We created the FeatureTracker challenge to probe how well observers could track objects as their appearances changed in familiar or unfamiliar ways. We did this by sampling the starting state of each object’s color and/or shape from distributions that we varied from training to test time. To elaborate, object appearances for training were sampled from one distribution (Fig. 4, red cube); then observers were tested on objects with appearances sampled in four different ways: (i) colors/shapes from the same distribution, (ii) colors from a different distribution but shapes from the same distribution, (iii) colors from the same distribution but shapes from a different distribution, or (iv) colors from a different distribution and shapes from a different distribution (Fig. 4 and SI A.7.2). We systematically evaluated the abilities of humans and machine vision systems to track objects with changing appearances by comparing their performances and decision strategies on each test set.
@@ Results
Figure 6: Neural synchrony causes the CV-RNN to act more like humans than any other model tested. Decision correlations of humans and models for each testing condition: features out-of-distribution conditions (a) and object evolution conditions (b), computed using the Error consistency measure (Geirhos et al., 2020b). We present the same videos to humans and models and compare their errors. Each dot represents the error consistency averaged across human subjects. A higher number represents similar errors between humans and the model. The grey box represents the human inter-subject agreement. (c) Visualization of the phases of the hidden state of the complex attention mechanisms of our CV-RNN. The model affects a phase value for the target which is different from the one of the distractors and the background, and remains consistent across frames.
@@
All of the models that we tested, with the exception of the two based on visual Transformers, rivaled or exceeded human performance on videos with objects that were sampled from the same distribution as models and humans were trained on (Fig. 5a, top-left). The models performed similarly when tested on videos of objects with similar colors as training but different shapes (Fig. 5a, bottom-left). However, model performance fell precipitously when object colors were sampled from a different distribution at test time (Fig. 5a, right). In both of these conditions (same shapes/different colors, different shapes/different colors), the CV-RNN performed significantly better than the other models, which were close to chance accuracy. To further probe the abilities of models and humans to solve FeatureTracker, we generated versions where either the shape, color, or shape and color of objects were fixed to a single value. We first calibrated the performance of models and a new set of human participants against the first experiment by testing both on FeatureTracker videos where the color and shapes of objects varied according to the training distribution. In this case, humans were only rivaled by the CV-RNN, InT RNN, and two of the 3D-CNNs (Fig. 5b, top-left). Only the CV-RNN’s performance fell within the human confidence interval on the remaining conditions, which consisted of objects with fixed shapes but varying colors (Fig. 5b, bottom-left), varying shapes but fixed colors (top-right), or fixed colors and shapes (bottom-right; see SI Figs. A.12,  A.13 and  A.14 for extended benchmarks). We also include additional experiments with non-static background (Fig. A.17) and textured objects (Fig.A.18), as well as control experiments on the ability of the models to generalize to in-distribution features but out-of-distribution feature trajectories or to occlusions of the target in Figs. A.15 and A.16, as well as objects vanishing or moving in and out of the frame in Fig A.19.
@@
To better understand how well our implementation of neural synchrony captures human strategies for object tracking, we next compared the errors it makes on FeatureTracker to humans. To test this, we turned to the method introduced by Geirhos et al. (2020a), which uses Cohen’s $\kappa$ to compute error correlations between observers while controlling for their task accuracy. For any given version of FeatureTracker, we then computed human-to-human (intersubject) error consistency as the average $\kappa$ between all combinations of humans and model-to-human error consistency as the average $\kappa$ between a model and all humans. CV-RNN errors were far more consistent with humans than other models (Fig. 6a). The CV-RNN’s error consistency fell within the human intersubject interval on 6/8 versions of FeatureTracker, and was most similar to humans in the remaining two versions of FeatureTracker (expanded results in SI A.13.2, Figs.A.20 and A.21). The InT RNN was more consistent than the CV-RNN on FeatureTracker videos that were sampled from the same distribution as models were trained on (Fig. 6a, top-left), but was otherwise far less aligned with humans than the CV-RNN. In contrast, the remaining DNNs tended to make errors on different videos than humans. CV-RNN’s ability to learn neural synchrony makes it more similar to human accuracy and decision-making on FeatureTracker than any other model we tested. To understand how the CV-RNN uses neural synchrony, we visualized the phase of the model’s complex-valued hidden state $\phi_{xy}$ over FeatureTracker videos. The model learned to use the phase of its hidden state to track the position of targets independently of their color or shape. When the model correctly classified FeatureTracker videos, the phase of its attention smoothly tracked the target object and inhibited occluding distractors. When the model was incorrect, its phase tag jumped from object to object as it searched for the target (more examples in SI A.8.3).